# Light4GS: Lightweight Compact 4D Gaussian Splatting Generation via Context Model

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as an efficient and high-fidelity paradigm for novel view synthesis. To adapt 3DGS for dynamic content, deformable 3DGS incorporates temporally deformable primitives with learnable latent embeddings to capture complex motions. Despite its impressive performance, the high-dimensional embeddings and vast number of primitives lead to substantial storage requirements. In this paper, we introduce a Lightweight 4DGS framework, called Light4GS, that employs significance pruning with a deep context model to provide a lightweight storage-efficient dynamic 3DGS representation. The proposed Light4GS is based on 4DGS that is a typical representation of deformable 3DGS. Specifically, our framework is built upon two core components: (1) a spatio-temporal significance pruning strategy that eliminates over 64% of the deformable primitives, followed by an entropy-constrained spherical harmonics compression applied to the remainder; and (2) a deep context model that integrates intra- and inter-prediction with hyperprior into a coarse-to-fine context structure to enable efficient multiscale latent embedding compression. Our approach achieves over 12× compression and increases rendering FPS up to 20% compared to the baseline 4DGS, and also superior to frame-wise state-of-the-art 3DGS compression methods, revealing the effectiveness of our Light4GS in terms of both intra- and inter-prediction methods without sacrificing rendering quality.
1 Introduction
Novel view synthesis (NVS) aims to to reconstruct 3D scenes from only a collection of multi-view images and has become a hot spot in computer vision. Traditional NVS approaches often fail to provide a high-fidelity experience in real-time, as they suffer from complex structures [39] or cumbersome rendering processes [28]. Recently, 3D Gaussian Splatting (3DGS) [16] has gained great attention as a new NVS paradigm that achieves both state-of-the-art (SOTA) rendering quality and speed. By representing scenes with anisotropic Gaussian primitives, 3DGS enhances spatial expressiveness and adaptively refines photorealistic details across different viewports. Furthermore, its efficient differentiable splatting pipeline facilitates both rapid and explicit training, making 3DGS highly adaptable for interactive 3D applications.
@@ 1 Introduction
Figure 1: Motivation of 4DGS compression. Deformable primitives and hexplanes account for more than 99% of 4DGS’s storage (top-right). More than 60% of primtives show almost 0 contribution to the final rendering and the hexplane exhibits strong inter-correlation (top-left). Considering these, we introduce STP to prune insignificant primitives and employ MHCM to compress hexplanes, with the three key technical components (bottom).
@@
With increasing demand for immersive experiences, extending 3DGS to dynamic scenes has become a growing research focus in this field. A straightforward approach involves training separate GS models for each timestamp, but this strategy quickly inflates storage, especially for long sequences. The time-variant 3DGS [21, 48] addresses per-frame training redundancy by modeling Gaussian attributes as time functions, albeit with significant training and storage costs. In contrast, the deformable 3DGS reduces modeling complexity by employing a small MLP to deform primitives trained from vanilla 3DGS. In this category, deformation of primitives is typically assisted by latent embeddings such as positional embeddings [49, 2], hash tables [45] and low-rank grids [43]. Among them, 4DGS [43] which utilizes multiscale hexplanes as latent embeddings is the most representative one. By mapping 4D space into six orthogonal planes, hexplane efficiently reduces computational complexity, making 4DGS more practical for real applications than the time-variant 3DGS. However, 4DGS’s storage requirements remain prohibitively high for real-world streaming. For instance, while current 5G networks stably support high-definition video at bitrates below 3 Mbps [33, 23], 4DGS requires 8–10 Mbps for comparable scenes. Fortunately, the two largest components of 4DGS show considerable redundancy and are prime candidates for compression. As illustrated in Fig. 1, over 60% of its primitives have minimal impact on rendering and can be pruned without loss in quality. Moreover, multiscale hexplanes exhibit strong correlations, suggesting that deep context modeling could greatly mitigate its information redundancy. Motivated by these insights, we aim to achieve effective 4DGS compression at a lightweight level.
In this paper, we propose Light4GS, a lightweight 4D Gaussian Splatting compression framework based on 4DGS [43], which leverages efficient pruning and deep context modeling. The motivation behind our approach is illustrated in Fig. 1. Our framework comprises two main modules: (1) a Spatio-Temporal significance Pruning method (STP) that reduces deformable primitives by evaluating their importance across different rendering views at all timestamps. STP not only saves storage but also streamlines training by automatically excluding gradients of both low-importance primitives and hexplane features. Since spherical harmonics (SH) coefficients account for over 80% of the primitive’s data volume, entropy-constrained compression is further applied to compress the SH; and (2) a Multiscale Hexplane Context Model (MHCM) that leverages intra- and inter-scale correlations to compress multiscale hexplanes: for the lowest-scale hexplane, a checkerboard context combined with either hyperprior or inter-plane context is employed to reduce spatio-temporal redundancies; high-scale hexplanes are compressed using inter-scale context in a coarse-to-fine fashion with reduced training complexity and improved compression effectiveness. In summary, our contributions are three-fold:
@@
  • We propose the first lightweight 4DGS compression framework that integrates primitive pruning with a deep context model for storage-efficient dynamic 3DGS.
  • We design a spatio-temporal significance pruning technique to prune redundant primitives and customized coarse-to-fine deep contexts to compress multiscale hexplanes. We also apply adaptive quantization during training and entropy constraints to compress SH.
  • Experiments on synthetic and real-world scenes show Light4GS achieves 10-200x compression over SOTA dynamic NVS methods without sacrificing quality.
2 Related Works
3DGS Compression: Recent studies in 3DGS compression can be categorized into either post-processing-based or learning-based approaches. Post-processing methods include significance pruning [9, 18], codebooks [18], knowledge distillation [9], and quantization [30] to either prune primitives or compress Gaussian attributes. Thanks to 3DGS’s explicit structure, point cloud compression techniques such as octree encoding [36, 15], graph signal processing [47] and region-adaptive hierarchical transform (RAHT) [44] have also been adapted for 3DGS compression with structural considerations. Learning-based methods encompass strategies like Scaffold [26], which optimize neural Gaussians with residual components to reduce local redundancy, as well as entropy-constrained approaches [8, 24] that leverage global context for compressing 3DGS more effectively.
Dynamic 3DGS: While 3DGS has been successfully compressed, addressing the complexity of time modeling in dynamic 3DGS presents another challenge. Time-variant methods either decompose 4D Gaussians into time-conditioned 3D Gaussians with auxiliary components [48] or represent positions and attributes as functions of time to capture object motions [21]. Conversely, time-invariant (deformation-based) methods learn implicit position mappings that deform primitives using latent embeddings such as positional embeddings [49, 2], hash tables [45], and low-rank grids [43]. While the latter category achieves superior quality-storage performance, it lacks compact optimization for 3DGS as reviewed earlier and does not compress the latent embeddings, which we explore in the following.
@@ 2 Related Works
Latent Embedding Compression: NVS methods utilizing latent embeddings [43, 38, 29, 11, 5] achieve high quality results, but also lead to increased storage overhead. To make embeddings smaller, [29] maps a large feature grid into a compact hash table, while [38] distills a high-dimensional feature grid into a binary grid. In dynamic NVS, storage overhead for latent embeddings become even more pronounced; low-rank approximations, such as wavelet decomposition [37], vector-matrix decomposition [6], and HexPlane projection [5, 11], provide promising solutions by mitigating dimensional redundancy. Moreover, contextual dependencies among neighboring elements provide further compression potential, as explored in image compression [13, 14], as well as in video compression [19]. To extend these advances into 3D representations, [7, 8] introduce spatial and hash-grid contexts to capture spatial redundancy in neural radiance field (NeRF) and 3DGS latent embeddings, which achieve nearly 100 compression gain over their vanilla implementations.
@@ 3 Preliminary
4 Method
We present a storage-efficient deformable 3DGS compression framework that combines STP to reduce deformable primitives with MHCM for compressing multiscale hexplanes. For further optimization, the framework employs entropy-constraint SH compression and adaptive quantization to tailor the quantization of different plane types. We begin with an introduction to our pruning strategy, i.e., STP, in Sec. 4.1, followed by a detailed description of our deep context model and its coding process, i.e., MHCM, in Sec. 4.2. The remaining sections address our optimization techniques and loss functions.
4.1 STP: Spatio-temporal Significance Pruning
@@
For this concern, a critical challenge in 4DGS is accurately capturing the global impact of each primitive on rendering quality. In 4DGS, each primitive is capable of deforming over time to account for scene dynamics, and this deformability enables us to assess the significance of one primitive across multiple frames. Building on this capability, we propose a spatio-temporal pruning strategy (STP) that evaluates each Gaussian’s contribution by calculating its intersection with viewing pixels over all timestamps. The global significance score of the $j$-th deformable primitive $\text{S}_{j}$ is defined by aggregating its contributions across all timestamps and views as:
$\text{S}_{j}=\sum_{t=1}^{\text{T}}\sum_{i=1}^{\text{MHW}}\mathbf{1}\left(\mathbf{G}+\Phi(f_{h},t)),\ r_{i,t}\right)\cdot\sigma_{j}\cdot\gamma(\Sigma_{j,t}),$ (6)
[…]
4.2 MHCM: Multiscale Hexplane compression with Context Model
@@
We use deep context model to reduce information uncertainty within multiscale hexplanes to achieve a minimized entropy loss. Specifically, each hexplane feature is modeled as a Gaussian distribution conditioned on a carefully designed deep context. This context-conditioned distribution is parameterized by a deep network (denoted as MLP() here for simplicity), which takes the context as input to predict each feature’s distribution parameters. These parameters are then used to compute the entropy of hexplane features, and subsequently encode features using Arithmetic Coding (AC) [17]. The total entropy is incorporated into the 4DGS optimization to balance the rate-distortion (RD) trade-off. In Light4GS, we propose a Multiscale Hexplane Context Model, MHCM that leverages intra- and inter-scale redundancies to compress multiscale hexplanes. Due to the mirroring nature of encoding and decoding, we present the context model from the encoding perspective, elaborating it in a coarse-to-fine manner in the following.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as an efficient and high-fidelity paradigm for novel view synthesis. To adapt 3DGS for dynamic content, deformable 3DGS incorporates temporally deformable primitives with learnable latent embeddings to capture complex motions. Despite its impressive performance, the high-dimensional embeddings and vast number of primitives lead to substantial storage requirements. In this paper, we introduce a Lightweight 4DGS framework, called Light4GS, that employs significance pruning with a deep context model to provide a lightweight storage-efficient dynamic 3DGS representation. The proposed Light4GS is based on 4DGS, which is a typical representation of deformable 3DGS. Specifically, our framework is built upon two core components: (1) a spatio-temporal significance pruning strategy that eliminates over 64% of the deformable primitives, followed by an entropy-constrained spherical harmonics compression applied to the remainder; and (2) a deep context model that integrates intra- and inter-prediction with hyperprior into a coarse-to-fine context structure, enabling efficient multiscale latent embedding compression. Our approach achieves over 12× compression and increases rendering FPS up to 20% compared to the baseline 4DGS, and also superior to frame-wise state-of-the-art 3DGS compression methods. Experiment results show the effectiveness of our Light4GS in terms of both intra- and inter-prediction methods without sacrificing rendering quality. The code is available at https://github.com/Evan-sudo/Light4GS.
I Introduction
@@
Figure 1: Motivation of 4DGS compression. Deformable primitives and hexplanes account for more than 99% of 4DGS’s storage (top-right). More than 60% of primitives show almost 0 contribution to the final rendering and the hexplane exhibits strong inter-correlation (top-left). Considering these, we introduce STP to prune insignificant primitives and employ MHCM to compress hexplanes, with the three key technical components (bottom).
Dynamic view synthesis has become a central problem in 3D vision, with the goal of generating photorealistic renderings of dynamic scenes from arbitrary viewpoints [66, 67, 11]. 3D Gaussian Splatting (3DGS) has recently gained significant attention for its ability to deliver faster rendering and higher visual quality than neural radiance field (NeRF)-based counterparts [4] in terms of static view synthesis. It represents 3D scenes using anisotropic Gaussian primitives and is trained end-to-end via a highly efficient differentiable rasterization pipeline. Similar to NeRF, 3DGS can be extended with a temporal dimension to support dynamic view synthesis.
@@
Recent efforts in dynamic 3DGS have led to two main categories of approaches: time-variant and deformation-based. Time-variant methods [30, 63] model Gaussian attributes as continuous functions of time, enabling shared representations across frames and avoiding per-frame training. Although impressive results are achieved, time-variant methods require substantial memory and incur considerable training overhead. Deformable 3DGS reduces modeling complexity by employing a lightweight multilayer perceptron (MLP) to deform primitives learned from a static 3DGS template. This idea is inspired by deformable NeRF approaches [46, 42], deformation of primitives is typically assisted by latent embeddings such as positional embeddings [64, 2, 23], hash tables [56] and low-rank grids [54]. Among them, 4DGS [54], which utilizes multiscale hexplanes as latent embeddings, is the most representative one. By mapping 4D space into six orthogonal planes, hexplane efficiently reduces computational complexity, making 4DGS more practical for real applications than the time-variant solution. Nonetheless, 4DGS’s storage requirements remain prohibitively high for real-world streaming. While current 5G networks stably support high-definition video at bitrates below 3 Mbps [44, 32], 4DGS requires 8–10 Mbps for comparable scenes. This gap shows that 4DGS is still too large for practical use in real-world streaming settings.
Fortunately, the two largest components of 4DGS show considerable redundancy and are prime candidates for compression. As illustrated in Fig. 1, over 60% of its primitives have minimal impact on rendering and can be pruned without loss in quality. Moreover, multiscale hexplanes exhibit strong correlations, suggesting that deep context modeling could greatly mitigate their information redundancy. Motivated by these insights, we aim to achieve effective dynamic 3DGS compression based on the 4DGS anchor.
In this paper, we propose Light4GS, a lightweight 4D Gaussian Splatting compression framework based on 4DGS [54], which leverages efficient pruning and deep context modeling. The motivation behind our approach is illustrated in Fig. 1. To improve spatio-temporal encoding, we first apply Principal Component Analysis (PCA) to the spatial coordinates of the Gaussian primitives, aligning them with the principal directions of the scene. Then, a non-linear contraction is applied to the spatial coordinates to address the unbounded nature of 4DGS, ensuring that queried points remain within the hexplane boundaries. Next, our framework compresses both the canonical 3DGS and its associated hexplane features: 1) For canonical 3DGS, we introduce a Spatio-Temporal Significance Pruning (STP) strategy to globally remove unimportant primitives. STP evaluates each primitive’s contribution across multiple views and timestamps, and prunes deformable primitives based on their significance ranking. This also accelerates training by skipping gradient updates for pruned primitives and their associated hexplane features. Since spherical harmonics (SH) coefficients account for over 80% of the primitive’s data volume, entropy-constrained compression is further applied to compress the SH. 2) For hexplane compression, we propose a Multiscale Hexplane Context Model (MHCM) by leveraging both intra-scale and inter-scale redundancies in a coarse-to-fine manner. At the lowest scale, we adopt a checkerboard context model combined with either a learned hyperprior or an inter-plane context, depending on the plane type. Space-time planes are encoded first using latent hyperpriors extracted from a deep encoder. The decoded results are then time-averaged to form an inter-plane context, which guides the compression of space-only planes without introducing extra side information. At higher scales, we introduce inter-scale context by decoding finer hexplanes using bilinearly interpolated features from the preceding coarser scale. This design captures cross-scale redundancy, improves compression efficiency, and reduces training complexity. In summary, our contributions are as follows:
@@
  • We propose Light4GS, a lightweight 4DGS compression framework that integrates primitive pruning with a deep context model for storage-efficient dynamic view synthesis.
  • We introduce a spatio-temporal significance pruning strategy to remove redundant primitives, and apply entropy-constrained optimization to compress the spherical harmonics (SH) coefficients of the remaining ones. We design a multiscale Hexplane Context Model (MHCM) that efficiently compresses hexplane features by leveraging both intra- and inter-scale redundancies in a coarse-to-fine manner with adaptive quantization.
[…]
II Related Works
II-A Dynamic view synthesis
Dynamic view synthesis aims to generate novel views of dynamic scenes from 2D images captured across different time instances, which remains a challenging problem [65, 20]. Early approaches extend NeRF [39] to 4D by either deforming a canonical neural radiance field or explicitly modeling it as time-varying [46, 42, 50, 16]. While effective, these methods suffer from slow rendering speeds due to their computationally expensive per-ray queries. To overcome this, recent work introduces dynamic 3DGS, which extends 3DGS into the temporal domain. These methods generally fall into two categories: time-variant and deformation-based. Time-variant methods encode temporal changes directly into Gaussian attributes using polynomial, Fourier, or learned functions [37, 30], or by explicitly extending 3D Gaussian primitives into 4D space [63]. Although these methods often achieve high reconstruction quality, they incur significant training and storage costs. In contrast, deformation-based methods [54, 35, 2] represent a scene using a canonical set of 3D Gaussians and a deformation field to capture temporal variations. While such deformation fields may struggle to perfectly track highly dynamic objects, they are more compact and easier to train. These deformation models are often enhanced with neural feature encodings, such as hash grids [10, 40], per-Gaussian embeddings [2], or low-rank grids [8, 7]. A representative example is 4DGS [54], which combines a canonical Gaussian set with a Hexplane-based deformation field, achieving competitive performance with relatively low storage on both multi-view and monocular video datasets. We adopt it as the anchor design for our 4DGS compression method.
II-B 3DGS and its Compression
3DGS has emerged as a powerful alternative for 3D scene representation endowed with learnable shape and anisotropic attributes. The implementation is efficient as it only requires hundreds of multiview photos for a complex scene. By adopting differentiable splatting and tile-based rasterization, 3D Gaussians are optimized during training to best fit their local 3D regions. Its explicit structure also leads to faster training and rendering speeds than SOTA NeRF algorithms. Recent studies in 3DGS compression can be categorized into various categories. Post-processing methods, generally applied after training, reduce Gaussian primitives and compress their attributes utilizing neural network-inspired compression. Key strategies include significance pruning [12, 27, 33], codebooks [27], knowledge distillation [12], and quantization [41, 61]. Thanks to the explicit structure of 3DGS, classical point cloud compression techniques such as octree encoding [47, 22], graph signal processing [60], and region-adaptive hierarchical transform (RAHT) [55] have been adapted for 3DGS compression with structural priors. A different line of research focuses on leveraging spatial relations to impose structured organization on otherwise unstructured Gaussians. These methods often collapse dimensions to arrange Gaussian primitives into regular 2D grids, facilitating efficient encoding. For example, SUNDAE [62] combines pruning strategies with spectral graph analysis to guide the creation of compact models. Mini-Splatting [13] proposes a spatially aware spawning mechanism to regulate the expansion of primitives. In contrast, Scaffold-GS [36] uses an anchor-based architecture, where Gaussian attributes are predicted via neural inference, reducing the number of parameters while improving reconstruction quality. Meanwhile, learning-based methods utilize VAE-style image compression frameworks that integrate hyperpriors and context modeling [10, 34] for more effective 3DGS compression.
II-C Latent Embedding Compression
@@
NVS methods utilizing latent embeddings [54, 49, 40, 15, 7] achieve high-quality results, but also lead to increased storage overhead. To make embeddings smaller, [40] maps a large feature grid into a compact hash table, while [49] distills a high-dimensional feature grid into a binary grid. In dynamic NVS, storage overhead for latent embeddings becomes even more pronounced; low-rank approximations, such as wavelet decomposition [48], vector-matrix decomposition [8], and HexPlane projection [7, 15], provide promising solutions by mitigating dimensional redundancy. Moreover, contextual dependencies among neighboring elements provide further compression potential, as explored in image compression [19, 18], as well as in video compression [28, 38]. To extend these advances into 3D representations, [9, 10] introduce spatial and hash-grid contexts to capture spatial redundancy in NeRF and 3DGS latent embeddings, which achieve nearly 100 compression gain over their vanilla implementations.
@@ III Preliminary
IV Method
We present Light4GS, a storage-efficient deformable 3DGS compression framework that combines STP to reduce deformable primitives with MHCM for compressing multiscale hexplanes, as shown in Fig. 2. We first present modifications to the vanilla 4DGS representation in Sec. IV-A. Then, we introduce our pruning strategy, i.e., STP, in Sec. IV-B, followed by a detailed description of our deep context model and its coding process, i.e., MHCM, in Sec. IV-C. The remaining sections address our optimization techniques and loss functions.
IV-A PCA-guided Hexplane Query for Unbounded 3DGS
Figure 3: Gaussian splatting is unbounded, leading to queries outside the hexplane. Misaligned hexplane query directions with principal directions further reduce the amount of information the hexplane can learn (left); Nonlinear contraction constrains all points within the hexplane bounds, while PCA aligns query directions with principal directions (right).
The vanilla 4DGS encodes spatio-temporal variations using six orthogonal planes spanning all pairwise combinations of $(x,y,z,t)$. Despite this compact representation, we observe two key limitations (see Fig. 3): (1) while Gaussian primitives may occupy an unbounded 3D space, the hexplane grid is typically initialized within a fixed bounding box derived from the SfM point cloud, leading to a mismatch and uneven coverage of spatio-temporal features; and (2) the canonical axes $(x,y,z)$ are not necessarily aligned with the true principal directions of the scene, so the resulting non-principal projections can hinder the hexplanes from learning more informative latent features.
Motivated by how NeRF handles its unbounded issues, we adopt the non-linear spatial contraction of [5]:
$\text{contract}(\boldsymbol{\mu})=\begin{cases}\boldsymbol{\mu},&\|\boldsymbol{\mu}\|\leq 1\\ \left(2-\frac{1}{\|\boldsymbol{\mu}\|}\right)\frac{\boldsymbol{\mu}}{\|\boldsymbol{\mu}\|},&\|\boldsymbol{\mu}\|>1,\end{cases}$ (6)
to map Gaussian centers $\boldsymbol{\mu}$ into a bounded spatial domain. Since the temporal dimension $t$ is already finite, no contraction is applied to it.
For the second issue, we perform PCA on the spatial coordinates $\boldsymbol{\mu}$ to align them with the principal directions, before applying the nonlinear contraction:
$\boldsymbol{\mu}^{\prime}=\text{contract}\left(\frac{(\mathbf{P}(\boldsymbol{\mu}-\boldsymbol{\mu}_{0})}{\boldsymbol{\sigma}_{0}}\right),$ (7)
where $\mathbf{P}\in\mathbb{R}^{3\times 3}$ is the PCA rotation matrix, and $\boldsymbol{\mu}_{0},\boldsymbol{\sigma}_{0}$ and denote the mean and the variance vectors along the principal axes.
Using the transformed coordinates $\boldsymbol{\mu}^{\prime}$, we perform bilinear interpolation over the learnable multiscale hexplanes $\mathbf{R}^{l}_{c}\in\mathbb{R}^{h\times lN_{1}\times lN_{2}}$, where $c\in\binom{\{1,2,3,4\}}{2}$ are PCA axis indices. The final hexplane feature $f_{h}$ is obtained by aggregating interpolated features across all planes and scales:
$f_{h}=\bigcup_{l}\prod_{c}\text{interp}\left(\mathbf{R}^{l}_{c},(\boldsymbol{\mu}^{\prime},t)\right),$ (8)
This replaces the original query procedure in Eq. (5).
IV-B STP: Spatio-temporal Significance Pruning
@@
For this concern, a critical challenge in 4DGS is accurately capturing the global impact of each primitive on rendering quality. In 4DGS, each primitive is capable of deforming over time to account for scene dynamics, and this deformability enables us to assess the significance of one primitive across multiple frames. Building on this capability, we propose STP that evaluates each Gaussian’s contribution by calculating its intersection with viewing pixels over all timestamps. The global significance score of the $j$-th deformable primitive $\text{S}_{j}$ is defined by aggregating its contributions across all timestamps and views as:
$\text{S}_{j}=\sum_{t=1}^{\text{T}}\sum_{i=1}^{\text{MHW}}\mathbf{1}\left(\mathbf{G}+\Phi(f_{h},t)),\ r_{i,t}\right)\cdot T_{j,i,t}\cdot\sigma_{j}\cdot\gamma(\Sigma_{j,t}),$ (9)
where M is number of views at each timestamp, $\text{H},\text{W}$ are height and width of the rendered view. $\mathbf{1}(\cdot)$ is an indicator function that returns $1$ if the deformed primitive intersects with the viewing ray $r_{i,t}$ at timestamp $t$, and 0 otherwise. $T_{j,i,t}=\prod_{k<j}(1-\alpha_{k,i,t})$ denotes the per-pixel transmittance in alpha compositing, where $\alpha_{k,i,t}$ is the per-pixel opacity contributed by primitive $k$ at pixel $(i,t)$. $\sigma_{j}$ is the opacity of the primitive, and $\gamma(\Sigma_{j,t})$ is the normalized volume of primitive $j$ calculated from its covariance matrix $\Sigma_{j,t}$ at time $t$. During training, we periodically rank deformable primitives according to their spatio-temporal significance scores, $\text{S}_{j}$, and prune the least significant ones based on a predefined ratio. This approach greatly improves training efficiency by reducing the number of primitives to be processed, while optimizing the remaining primitives. Consequently, the retained primitives adapt to mitigate any potential quality loss from STP, preserving rendering quality despite the reduced model complexity.
IV-C MHCM: Multiscale Hexplane compression with Context Model
@@
We use a deep context model to reduce information uncertainty within multiscale hexplanes to achieve a minimized entropy loss. Specifically, each hexplane feature is modeled as a Gaussian distribution conditioned on a carefully designed deep context. This context-conditioned distribution is parameterized by a deep network (denoted as MLP() here for simplicity), which takes the context as input to predict each feature’s distribution parameters. These parameters are then used to compute the entropy of hexplane features, and subsequently encode features using Arithmetic Coding (AC) [26]. The total entropy is incorporated into the 4DGS optimization to balance the rate-distortion (RD) trade-off. In Light4GS, we propose MHCM, which leverages intra- and inter-scale redundancies to compress multiscale hexplanes. Due to the mirroring nature of encoding and decoding, we present the context model from the encoding perspective, elaborating it in a coarse-to-fine manner.
IV-C1 Lowest-Scale Hexplane Compression